the half-life of 234u, uranium-234, is 2.52 105 yr. if 98.4% of the uranium in the original sample is present, what length of time (to the nearest thousand years) has elapsed?
Answer by stanbon(57377)

the half-life of 234u, uranium-234, is 252,105 yr. if 98.4% of the uranium in the original sample is present, what length of time (to the nearest thousand years) has elapsed?
-----
A(t) = Ao*(1/2)^t
-----
Solve:
0.984Ao = Ao(1/2)^(t/252,105)
-----
(1/2)^t/252,105 = 0.984
---
Take the log of both sides:
(t/252,105) = log(0.984)/log(1/2) = 0.0233
---
t = 252,105*0.0233 = 5866.43 years
=======================================
Cheers,
stan H.